OpenStudy (angel_kitty12):

URGENT: can someone help me solve this ASAP The cube root of 4 times the sixth root of 8 3sqrt4•6sqrt8 I'm really confused here. Thanks

OpenStudy (malcolmmcswain):

Using a calculator, the cube root of 4 is approximately 1.58740105197.

OpenStudy (malcolmmcswain):

Also, the sixth root of 8 is approximately 1.41421356237.

OpenStudy (malcolmmcswain):

Multiply: 1.58740105197*1.41421356237=2.2449240966 So, the cube root of 4 times the sixth root of 8 is approximately equal to 2.2449240966

OpenStudy (anonymous):

\[4^{\frac{ 1 }{ 3 }}\times8^{\frac{ 1 }{ 6 }}\]

OpenStudy (angel_kitty12):

I turned them into fraction format and got 32^5/6 but do I simplify it anymore?

OpenStudy (malcolmmcswain):

Oh, I thought you wanted this in decimal format.

OpenStudy (anonymous):

How did you get 5/6?

OpenStudy (angel_kitty12):

I'm not sure what to do after this step.

OpenStudy (anonymous):

So it's easier if you set both 4 and 8 to 2 to the power of something

OpenStudy (angel_kitty12):

I created a common denominator 4^2/3 times 6 and 8^1/6 times 3 to get the final as 32^ 5/6

OpenStudy (angel_kitty12):

After simplified

OpenStudy (anonymous):

\[4=2^{2} and 8=2^{3}\]

OpenStudy (angel_kitty12):

Wait I carried the wrong number. It's supposed to be 1/3 not 2/3

OpenStudy (anonymous):

Using that you can have a common value that is being raised to some exponent

OpenStudy (angel_kitty12):

Okay the exponents seem understandable but what about the fraction forms? Do you add that to the exponent?

OpenStudy (anonymous):

So \[4^{\frac{ 1 }{ 3 }}=2^{\frac{ 2 }{ 3 }}\] and \[8^{\frac{ 1 }{ 6 }}=2^{\frac{ 3 }{ 6}}\]

OpenStudy (anonymous):

So after you change the fraction so that you have a common denominator, you should be getting \[2^{\frac{ 7 }{ 6 }}\]

OpenStudy (angel_kitty12):

Oh okay I understand

OpenStudy (angel_kitty12):

And the answer should be 2 times the sixth root of 2 2• 6sqrt2

OpenStudy (angel_kitty12):

Wait question

OpenStudy (anonymous):

Technically that's correct but that wouldn't be the most simplified form

OpenStudy (angel_kitty12):

Why don't you multiply the 2 to get 4?

OpenStudy (anonymous):

What's the question?

OpenStudy (angel_kitty12):

4^7/6

OpenStudy (anonymous):

It's an exponent rule Well you can solve it out for proof as to why you can't just multiply the 2's to get 4 Let's pretend that your exponent isn't a fraction but a whole number (since it's easier to understand) EXAMPLE \[2^{3}\times2^{2}=8\times4=32\] If you had multiplied the 2's together and said that \[2^{3}\times2^{2}=4^{5}\] that would not be equal to 32 (which is the correct answer)

OpenStudy (anonymous):

And 2^5=32

OpenStudy (angel_kitty12):

Oh alright. Thank you so much for being patient with me and helping me. I understand everything now. Thank you!!!

OpenStudy (anonymous):

np :)